The product of 7 and the square of a number is 112. Find the number.
Answer by mananth(16946)

The product of 7 and the square of a number is 112. Find the number.
let the number be x
square wll be x^2
7x^2=112
7x^2= 112
x^2 = 112/7
x^2=16
x=4